OpenStudy (anonymous):

[3.07] Choose the correct procedure for calculating (f − g)(−2) if f(x) = 13x2 and g(x) = 7x − 5.

OpenStudy (anonymous):

find f(x) - g(x) and plug in -2

OpenStudy (shaik0124):

(f-g)(-2)=f(-2)-g(-2) =13(-2)^2-7(-2)-5 =52+14-5 =61

OpenStudy (anonymous):

Thanks and all you did was change the x's to 2